OpenStudy (anonymous):

ln(y)=21ln(x)+ln(10). Express y in terms of X

OpenStudy (raden):

21ln(x) = ln(x)^21 next simplify right side. ln(x)^21 + ln(10) = ln((x^21)10)

OpenStudy (raden):

so, ln(y) = ln((x^21)10) know, can u get value of y in terms of x?

OpenStudy (anonymous):

I have no idea :(

OpenStudy (raden):

u can cancel out of ln